P346

SPECTROSCOPY OF THE PRODUCTS OF THE REACTIONS BETWEEN ABLATED TRANSITION MET-ALS AND SMALL HYDROCARBONS

\underline{JOHN A. DICKINSON} AND TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

The products of the reactions between ablated transition metals and small hydrocarbons have been investigated using laser induced fluorescence. In the low resolution (FWHM=0.05cm^{−1}) spectrum of the products of ablated Pt with C_{3}H_{3} or CH_{4} a series of weak features close to the ^{1}Π ← ^{1}Σ^{+} band of PtC^{a} have been identified. These features are broad and exhibit no reso lveable rotational structure. All the bands disappear when the C_{3}H_{3} or CH_{4} is replaced with NH_{3}, H_{2} or CS_{2}. The dispersed fluorescence spectra of these bands displays a low frequencey progression (ω_{e} ≈ 250cm^{−1}) close to that observed in Pt_{2}, however, they are not Pt_{2}^{b}^{c}. On this basis, these features are tentatively assigned to Pt_{2}CH_{x}. Progress on the analysis of these bands, higher resolution work and new work on Fe and other transition metals with small hydrocarbons will be presented.

aT. C. Steimle, K. Y. Jung and B.-Z. Li, J. Chem. Phys., 102, 5937, (1995)

bM. Morse, Private communication

cS. Taylor, G. W. Lemire, Y. M. Hamrick, Z. Fu and M. D. Morse, J. Chem. Phys., 89, 5517, (1988)